OpenStudy (anonymous):

What is the formula for calculating v/v % ? Indicate what each variable represents

OpenStudy (chmvijay):

Volume by volume % is nothing but = volume of solute / total volume of solution * 100 example if u add 5 ml of alcohol to 90 ml of water then the V/V % of alcohol will be = 5 /90+5 *100 =5/95 *100 = 5.26 %